Email account password consists of five characters: one lowercase letter and four digits 0 – 9

  1. What is a percentage of passwords with repeated digits?
Answer #1

Number of passwords possible = 26*10*10*10*10 = 260000

Number of passwords with unique digits = 26*10*9*8*7 = 131040

Hence,

Percentage of passwords with unique digits = (131040/260000)*100 = 50.4%

Therefore,

Percentage of passwords with repeated digits = 100% - 50.4% = 49.6%